area of square is 400. find the cost of boundary wall if the rate of per meter is rs 5.
Answer by mananth(12270)

area of square =400
Area = L^2 L = length of side
L^2=400
L= sqrt 400
L=20
..
Perimeter of square = 4*L
=4*20=80
Rate /meter = Rs.5
For 80 meters it will be 5*80=Rs400